If y varies directly as x and y=12 when x=5, find the equation of variation.
Answer by Alan3354(69443)

If y varies directly as x and y=12 when x=5, find the equation of variation.
-------------------
y = kx where k is a constant
12 = k*5
k = 12/5 = 2.4